### v1

# Measuring Psychological Depth in Language Models

Abstract
Evaluations of creative stories generated by large language models (LLMs) often focus on objective properties of the text, such as its style, coherence, and toxicity. While these metrics are indispensable, they do not speak to a story’s subjective, psychological impact from a reader’s perspective. We introduce the Psychological Depth Scale (PDS), a novel framework rooted in literary theory that measures an LLM’s ability to produce authentic and narratively complex stories that provoke emotion, empathy, and engagement. We empirically validate our framework by showing that humans can consistently evaluate stories based on PDS ( Krippendorff’s alpha). We also explore techniques for automating the PDS to easily scale future analyses. GPT-4o, combined with a novel Mixture-of-Personas (MoP) prompting strategy, achieves an average Spearman correlation of with human judgment while Llama-3-70B scores as high as for empathy. Finally, we compared the depth of stories authored by both humans and LLMs. Surprisingly, GPT-4 stories either surpassed or were statistically indistinguishable from highly-rated human-written stories sourced from Reddit. By shifting the focus from text to reader, the Psychological Depth Scale is a validated, automated, and systematic means of measuring the capacity of LLMs to connect with humans through the stories they tell.
[…]
Measuring Psychological Depth in Language Models
Fabrice Harel-Canada Hanyu Zhou Sreya Mupalla Zeynep Yildiz Amit Sahai Nanyun Peng University of California, Los Angeles fabricehc@cs.ucla.edu
1 Introduction
Stories play a crucial role in our understanding of ourselves and the world around us Langer (1942); Kaniss (1991). As large language models (LLMs) are increasingly deployed in narrative design and creation, their growing impact on how stories are told calls for a deeper understanding of their narrative power. Current evaluations of LLM-generated stories often focus on objective properties of the text such as discourse structure Liu et al. (2024), fluency Gómez-Rodríguez and Williams (2023), style Wang et al. (2024b), creativity Chakrabarty et al. (2023), diversity Gehrmann et al. (2021), toxicity and bias Wang et al. (2024a). However, it is crucial to extend these evaluations to accommodate the subjective, psychological impact stories have on readers. While some studies have recently explored aspects like empathy Ma et al. (2020); Sabour et al. (2022) and engagement Ghazarian et al. (2020); Xu et al. (2022), they do not fully capture the multifaceted and interconnected nature of the reading experience.
Recognizing this gap, our study introduces a novel approach to measuring the psychological depth of short stories. We present the Psychological Depth Scale (PDS), drawing inspiration from two literary theory frameworks: reader-response criticism Holland (1989) and text world theory Gavins (2007a). Reader-response criticism emphasizes the reader’s subjective experience, while text world theory examines how readers cognitively construct a nuanced and realistic mental model of a story. Using related search terms, we conducted an extensive literature review of 95 peer-reviewed articles and books, identified 143 different evaluation criteria, and merged many of the broader themes into five key metacomponents: empathy, engagement, emotion provocation, authenticity, and narrative complexity. By recognizing the dual roles of authors and readers in creating and interpreting narratives, we aim to offer a more streamlined and comprehensive framework for assessing the psychological depth of creative content (Figure 1).
To empirically validate the PDS, we conducted a study involving non-expert humans, who are increasingly engaging with creative content generated by LLMs. We enlisted five undergraduate students from a US public university to analyze a dataset of 97 stories authored by humans and five contemporary LLMs. The raters provided psychological depth ratings and predictions on human or machine authorship, with detailed justifications. This approach addresses three critical research questions to explore the intersection of LLMs and the psychological nuances of creative writing.
RQ1.
How consistently can well-informed humans judge psychological depth? The Psychological Depth Scale achieved an average Krippendorff’s alpha of , indicating a significant level of agreement among raters and affirming its validity as a reliable instrument for assessing fictional short stories.
RQ2.
To what extent can psychological depth be measured automatically? Leveraging our novel Mixture-of-Personas prompting strategy, GPT-4o achieved a Spearman correlation of with human judgment while Llama-3-70B attains correlations as high as for empathy and for narrative complexity. These results highlight that while no single LLM excels at predicting all components of psychological depth, a strategic combination of different LLMs shows significant promise for automating PDS analyses.
RQ3.
How do stories written by amateur humans and LLMs manifest psychological depth? Starting from the same Reddit premise, stories generated by GPT-4 surpassed popular human-authored stories with statistical significance on narrative complexity and empathy while being statistically indistinguishable on all other components. This constitutes a notable progression in the capacity of some LLMs for deep and impactful storytelling.
Overall, our findings validate the Psychological Depth Scale as an effective, automated, and systematic means of measuring how well LLMs connect with humans through storytelling. Remarkably, our results reveal that GPT-4 already matches or exceeds the quality of respected stories from Reddit, with of readers believing GPT-4’s stories to be human-written.111Code and dataset will be released upon acceptance.
2 The Psychological Depth Scale
The Psychological Depth Scale (PDS) aims to comprehensively assess the psychological depth of human and machine-authored narratives. PDS is underpinned by two reader-centered theoretical frameworks, reader-response criticism and text world theory, briefly summarized as follows. Reader-response literary theory centers the role and experience of the reader in narrative analysis. Instead of “what does this sentence mean?”, it asks “what does this sentence do?”, emphasizing readers’ role in the production of literary meaning Fish (1970); Mailloux (1976); Babaee and Montashery (2012). On a more cognitive-structural level, text-world theory suggests that people understand narratives by constructing "text worlds" Gavins (2007b); Canning (2017). Text-worlds are mental representations of a narrative dynamically evaluated and updated throughout the reading process Gavins (2007b). The practical advantages of both frameworks are observed across education Woodruff and Griffin (2017); Kunjanman and Abdul Aziz (2021), translation studies Tian and Wang (2019); Chan (2016), and consumer research Scott (1994); Kushneruk (2017).
The metrics for PDS are derived from an extensive literature review within cognitive psychology, media studies, and narrative analysis. 95 peer-reviewed articles and books were included in the final review. 143 evaluative components were extracted from the included works. We employed thematic analysis to group these candidates under five metrics: emotion provocation, empathy, engagement, authenticity, and narrative complexity. Below we analyze these metrics and discuss their impact. We note that while factors contributing to each metric may be complex, measuring their narrative achievement is relatively simple. As PDS is a reader-centered assessment tool, each metric is such that readers will know when a story achieves it. Our surveys reflect this fact.
Emotion Provocation (PROV) measures the narrative’s ability to elicit strong emotional responses, positive or negative. Theories of emotions range across neuroscience Clark-Polner et al. (2016); Mendes (2016), cognitive psychology Ortony et al. (1988); Lazarus (1991), and philosophy Sartre (1971); Nussbaum (2004). Contemporary approaches describe emotions as mental states marked by valence (positivity or negativity) and arousal (level of intensity) Barrett et al. (2016). Narrative evocation of emotions foster attention and interest, and evoked emotions may persist or recur after reading Mar et al. (2011).
Recent fMRI research shows that congruent (i.e. positive valence - high intensity, negative valence - low intensity) textual emotive expressions are more cognitively effective than conflicting (i.e. positive valence - low intensity, negative valence - high intensity) ones Citron et al. (2014); Megalakaki et al. (2019). The disparity between congruent and conflicting narrative emotions may be a factor in "compassion fatigue" Maier (2015); Kinnick et al. (1996) and "psychic numbing" Maier et al. (2016); Slovic (2007) exhibited by readers towards news about mass tragedies. Enhancing a story’s emotional impact is therefore more complex than increasing the amount and intensity of emotion-laden content as this may elicit weaker emotional responses. Given its impact and complexity, we contend that emotional provocation is an achievement of psychologically deep stories and is thereby a crucial metric in narrative quality assessment.
Empathy (EMP) captures narrative evocation of empathetic responses in readers, such as immersive identification with characters and cognitively partaking in narrated experiences Miall and Kuiken (2001); Oatley (2002); Zaki and Ochsner (2012). The neural and cognitive structure of empathy and its social and psychological impact is well-documented in the literature Davis (1994); Hoffman (1991); Uddin et al. (2007). Both behavioral studies and fMRI research demonstrate that empathetic identification facilitates a cognitive deployment shift in readers, privileging fictional perspective over one’s own perspective Kaufman and Libby (2012); Speer et al. (2009). Empathetic responses catalyze introspection and perspective-taking, facilitating prosocial behavior Grant (2008), emotional intelligence Mar et al. (2006), and insight into the human experience Davis et al. (1980); Gerdes et al. (2010); Singer et al. (2006). Empathetic narratives embody “suggestion structure[s]” that conjure up themes of shared human experiences through the use of tropes such as metaphor and metonymy Oatley (2002); Johnson-Laird and Oatley (2022). Narrative evocation of empathy is therefore a plausible indicator of the psychological depth of stories.
Engagement (ENG) assesses the ability of a story to captivate and maintain the reader’s attentional focus Busselle and Bilandzic (2009). Narrative engagement is a major component of transportation Gerrig (1993), whose positive social and cognitive effects have been extensively demonstrated Johnson (2012); Green and Brock (2000). As transportation consists of engagement and emotional response Green et al. (2008), we have chosen to treat each component separately. In fictional settings, engagement facilitates persuasion and strong attitudes, heavily influencing how much fun is experienced Goffman (1961). More engaged readers are more likely to lose track of time and fail to notice changes in their surroundings Busselle and Bilandzic (2008). In contrast, unengaged readers are more likely to be distracted or frustrated Bal and Veltkamp (2013). These results reveal the reciprocal role of engagement as both a predictor and an outcome of other narrative achievements.
Authenticity (AUTH) captures narrative expressions of genuine human experiences and emotions. Psychologically deep stories convey authentic aspects of human existence the reader can resonate with, and “feel one’s way in” (Einfühlung) Forster (2022) even when depictions involve radical mental and material differences. In philosophy, authentic expressions are considered to manifest Rousseau (2016) and construct Kierkegaard (1980) one’s true self, and capture the essence of human existence Heidegger (1962); Berlin (2000); Taylor (1991). Empirical research suggests that positive authenticity judgments help fulfill a social “need to belong” Newman and Smith (2016). Narrative realism enhances persuasive impact Petraglia (2009); Zwarun and Hall (2012) and promotes more interest in the story Green (2004); Hall (2003). Creating psychologically deep stories therefore involves ensuring that stories feel reflective of real-life complexities.
Narrative Complexity (NCOM) refers to the presence of rich and intricate storylines and character development, especially those that engender puzzled intrigue from the reader Somasundaran et al. (2018). Narrative complexity is often achieved through creative techniques such as nonlinear narration, plot twists, and double perspectivation Kiss and Willemsen (2017). Complex narratives present immersive puzzles that motivate readers to undertake mental restructuring and retroactive revision Kiss and Willemsen (2017); Cutting (2019). Exerting cognitive effort facilitates reader attention and interest Empson (1947); Steiner (1978), and can result in a more enjoyable reading experience Zunshine (2006). Narrative complexity is not narrational complexity: simple stories can be obscurely narrated, while complex stories can be free of complex language Cutting (2019). Furthermore, the extent to which narrative complexity is intersubjective depends on situational similarities and contextual cues Tikka and Kaipainen (2017). Thus, narrative complexity is a dynamic and reciprocal measure that goes beyond stylistic and structural choices.
3 PsychDepth Dataset
We developed a dataset to analyze the psychological depth of creative short stories, consisting of premise-response pairs in English with authorship metadata. The full dataset includes 495 stories: 45 human-written and 450 LLM-generated, each averaging about 450 words. Due to the time and cost involved in manual annotation, we also created a smaller subset of 97 stories using a stratified sampling method to balance prompt premises, authorship, and generation strategies. We used this smaller dataset in our Human Study in Section 4.
3.1 Human Stories
We collected human-authored stories from Reddit’s r/WritingPrompts,222https://www.reddit.com/r/WritingPrompts a popular online community with over 18 million users. This forum was chosen for its accessibility, thematic diversity, and the structured nature of its writing prompts. Aspiring writers respond to these prompts (called "premises") with their stories, which readers can upvote or downvote. While these stories may not represent the pinnacle of human writing, Reddit’s voting system allows us to approximate where LLM performance stands relative to variances in human writing quality. We categorized the stories based on their ranking position: Human-Advanced for top-voted stories, Human-Intermediate for medium-voted stories, and Human-Novice for low-voted stories, with average upvotes of 1434, 263, and 9, respectively. To ensure that quality was the primary factor influencing votes, we only selected stories that were posted within 24 hours of one another.
3.2 LLM Stories
When generating stories with LLMs, we employed a multifaceted approach involving five models, two prompting strategies, and three sampled generations. We intentionally restricted the model architecture to examine the impact of model size on psychological depth. In particular, we chose the Llama-2 family Touvron et al. (2023) with a variety of sizes (7B, 13B, and 70B). We added Vicuna-33B Chiang et al. (2023), which is fine-tuned based on Llama, to fill the size gap between 13B and 70B. We also included GPT-4 OpenAI et al. (2023) as the highest-performing LLM at the time. Appendix A.1 describes the quality control measures applied to LLM-generated stories.
3.2.1 Prompting Strategies
After extensive internal experimentation and prompt engineering, we developed two distinct prompting strategies to prime the LLMs for generating stories with exceptional psychological depth.
WriterProfile (WP): Prior work has shown that in-context impersonation of domain experts can improve LLM performance Salewski et al. (2023). Adopting this approach, we crafted a profile of a seasoned writer known for psychologically deep, engaging stories. This profile is prepended to the prompt to prime the LLM for exploring complex psychological states and evoking strong emotions.
Plan+Write (P+W): Inspired by prior work Yao et al. (2018); Goldfarb-Tarrant et al. (2019); Yang et al. (2022), the Plan+Write approach splits the writing process into two phases: Character Portraits and Story Composition. The Character Portraits phase augments a story prompt with details about the main characters, such as their emotional states and inner thoughts. The Story Composition phase expands on the premise and character profiles to produce the final story. Although other story components like setting, plot, and outline can be included, we found that adding multiple phases harmed the coherency and consistency of short stories. Therefore, we focused solely on character portraits in the Plan+Write approach.
4 Human Study
Participant Recruiting. We recruited undergraduate students in English and Psychology from a US public university using targeted fliers and emails. We hypothesized that participants with some aptitude for literary and psychological analysis would provide more valuable insights than a random cross-section of the population. From 47 applications, we selected the 5 most promising candidates based on their interests and previous experience in narrative and psychological analysis. Our goal was to engage informed laypeople with relevant backgrounds, bridging the gap between typical Amazon Mechanical Turkers and expert professionals.
Evaluation Protocol. We held an initial meeting to introduce the PDS and outline the annotation task. We provided a tutorial annotation session, followed up with short questions designed to help participants calibrate their understanding of each component of the PDS. We show our annotation instructions in Appendix A.5. After confirming task comprehension, participants were instructed to complete the annotations remotely within a week. To prevent annotator fatigue and promote careful annotation, stories were divided into batches of 20. The evaluation criteria required participants to (1) read the prompts and stories thoroughly; (2) rate the five components of psychological depth on a Likert scale from 1 to 5; (3) assess the likelihood of authorship on a Likert scale from 1 (LLM) to 5 (human); (4) provide explanations for ratings (optional).
Each story is evaluated by all 5 annotators. On average, annotations took approximately 7.8 hours to complete, and participants were compensated $100 each for their contributions. The study produced a rich dataset with 2,425 ratings for psychological depth, 485 authorship likelihood ratings, and 1,128 free-form justifications.
5 Results
5.1 RQ1. Consistency of Human Judgments
| Component | AUTH | EMP | ENG | PROV | NCOM | AVG |
| Llama-3-8B | 0.90 | 0.96 | 0.95 | 0.96 | 0.94 | 0.94 |
| Llama-3-70B | 0.93 | 0.96 | 0.91 | 0.95 | 0.96 | 0.94 |
| GPT-3.5 | 0.88 | 0.93 | 0.88 | 0.89 | 0.93 | 0.90 |
| GPT-4o | 0.92 | 0.92 | 0.90 | 0.95 | 0.92 | 0.92 |
| Human | 0.71 | 0.74 | 0.70 | 0.71 | 0.74 | 0.72 |
This question establishes the degree to which PDS is operational as a coherent framework for evaluating short stories. We employ the widely used Krippendorf’s alpha (K-) Krippendorff (2011) parameterized with an ordinal kernel metric to measure agreement among study participants’ Likert ratings. Human ratings exhibit notable consistency across the five components of psychological depth: Authenticity (), Empathy (), Engagement (), Emotion Provocation (), and Narrative Complexity () (See Human row of Table 1). Evaluator consistency establishes the practical potential of operationalizing PDS in literary studies and creative writing pedagogy.
5.2 RQ2. LLM-as-Judge for Measuring Psychological Depth
Acquiring human annotations is often costly and time-consuming. Recent work has demonstrated the potential of leveraging LLMs for automated evaluation of text summarization Liu et al. (2023) and creative generation Rajani et al. (2023). Correspondingly, we designed an automated evaluation procedure to study the degree to which four contemporary LLMs of various sizes and inference costs – Llama-3-8B, Llama-3-70B, GPT-3.5 (gpt-3.5-turbo-0125), and GPT-4o (gpt-4o-2024-05-13) – can assess psychological depth in a zero-shot fashion. For each story, we prompted LLMs with instructions similar to those provided to human participants but additionally required explanations that contextualize each numerical rating.
Exploring further benefits of in-context impersonation Salewski et al. (2023), we experimented with two different prompt settings: (1) a vanilla zero-shot baseline and (2) a novel Mixture-of-Personas (MoP) approach. For MoP, we queried GPT-4o to provide a set of relevant personas based on a description of the PDS components and task setting. We then repeated the zero-shot annotation with different personas designed to prime the LLMs for taking diverse perspectives towards textual analysis (e.g. see Table 9 in Appendix A.7). Similar to RQ1., we calculated Krippendorf’s alpha upon this set of ratings to measure the agreement among persona judgments. Before calculating Spearman Rank correlations between human and LLM judgments of psychological depth, all ratings were aggregated by a simple average to yield equivalently sized sets of consensus labels.
| Judge | AUTH | EMP | ENG | PROV | NCOM | Average | +MoP % |
| Llama-3-8B | 0.0786 | 0.4248 | 0.1981 | 0.3316 | 0.4641 | 0.2994 | – |
| Llama-3-8BMoP | 0.3175 | 0.4669 | 0.2272 | 0.3959 | 0.4665 | 0.3748 | 25.16% |
| Llama-3-70B | 0.2205 | 0.5790 | 0.2477 | 0.5181 | 0.5881 | 0.4307 | – |
| Llama-3-70BMoP | 0.2525 | 0.6793 | 0.2775 | 0.5695 | 0.6163 | 0.4790 | 11.23% |
| GPT-3.5 | 0.3867 | 0.4637 | 0.1800 | 0.3551 | 0.3289 | 0.3429 | – |
| GPT-3.5MoP | 0.4729 | 0.6024 | 0.1470 | 0.4182 | 0.5269 | 0.4335 | 26.43% |
| GPT-4o | 0.4537 | 0.5121 | 0.2923 | 0.4429 | 0.3840 | 0.4170 | – |
| GPT-4oMoP | 0.4820 | 0.6417 | 0.4218 | 0.5661 | 0.4241 | 0.5071 | 21.62% |
| +MoP % | 33.81% | 20.74% | 16.93% | 18.34% | 15.22% | 20.43% | – |
Table 1 shows the K- values, which indicate that the personas do inject some diversity of opinion into the ratings, though comparatively less than human participants. Remarkably, leveraging an ensemble of relevant personas increased correlation with human judgment relative to the vanilla zero-shot baseline. In Table 2, we present the zero-shot correlations between LLM-as-Judge and human evaluations for each PDS component. The data indicate that the Mixture-of-Personas significantly improves the correlation with human judgments across all models on average. For instance, the Llama-3-8B and GPT-4o models showed an average correlation improvement of and , respectively.
Among individual PDS components, authenticity and empathy show the most significant improvements. For instance, authenticity correlations for Llama-3-8B improve from to , a remarkable increase for that LLM and an average improvement of across all models. Empathy correlations for Llama-3-70B improve from to , the highest observed correlation across all components and models. These results suggest that MoP particularly enhances the models’ ability to judge certain components in ways that align more closely with human evaluations.
It is worth noting that no single model consistently outperforms the others across all components. While GPT-4o had the highest overall correlation of and excelled in quantifying authenticity and engagement, Llama-3-70B showed the best performance for measuring empathy (), narrative complexity (), and emotional provocation (). This variability underscores the importance of selecting and possibly combining multiple models depending on the specific evaluative criteria being prioritized.
Overall, the average percentage increase due to Mixture-of-Personas prompting across all models is approximately . These results demonstrate that diverse LLM opinions can more accurately reflect the multifaceted nature of human judgment, proving potential for more nuanced and human-like assessments by AI systems.
5.3 RQ3. Comparing Psychological Depth in Human and LLM Stories
To compare human- and LLM-written stories, we aggregated participant ratings by author and present the means and standard deviations in Table 3. We also computed statistical significances via pairwise t-tests between each combination of authors. Full results for that analysis can be found in Appendix A.8.
Remarkably, GPT-4 scored the highest on four out of five components of psychological depth, though only with statistical significance on empathy and narrative complexity. On authenticity, engagement, and emotional provocation, GPT-4 stories were statistically indistinguishable from stories by both advanced and intermediate human writers on Reddit. The table also illustrates a notable variance in scores across different levels of human writing where stories generated by Llama-2-7B are most comparable to those written by Human-Novice while GPT-4 is most similar to Human-Advanced. The smaller standard deviations also highlight GPT-4 as one of the most consistent authors in the study. Visualizations of these results can be found in Appendix A.6.
| Author | AUTH | EMP | ENG | PROV | NCOM | HUM |
| Llama-2-7B | 2.92 ± 1.23 | 2.62 ± 1.28 | 2.77 ± 1.27 | 2.64 ± 1.28 | 2.48 ± 1.30 | 2.87 ± 1.57 |
| Llama-2-13B | 2.96 ± 1.26 | 2.73 ± 1.23 | 2.51 ± 1.32 | 2.53 ± 1.33 | 2.43 ± 1.12 | 2.40 ± 1.58 |
| Vicuna-33B | 2.76 ± 1.37 | 2.59 ± 1.41 | 2.61 ± 1.52 | 2.59 ± 1.42 | 2.55 ± 1.32 | 2.44 ± 1.57 |
| Llama-2-70B | 3.09 ± 1.26 | 2.99 ± 1.23 | 3.01 ± 1.37 | 2.94 ± 1.26 | 2.73 ± 1.28 | 2.69 ± 1.56 |
[…]
| Human-Novice | 2.73 ± 1.22 | 2.07 ± 1.16 | 3.27 ± 1.39 | 2.67 ± 1.23 | 2.20 ± 1.32 | 3.93 ± 1.33 |
[…]
Beyond the five PDS components, participants were tasked with estimating authorship sources. Stories penned by humans averaged a rating of , compared to for the most "human-like" of LLMs, GPT-4. This discrepancy underscores the exceptional depth scores, particularly in light of the slightly higher suspicion of LLM authorship.
6 Discussion
Human vs. LLM Authorship Identification. On average, participants identified human vs. LLM authorship with only accuracy. For stories generated by GPT-4, accuracy dropped significantly to . Conversely, GPT-4’s accuracy in identifying authorship was , underscoring the challenge even for LLMs to distinguish between human and machine-generated content.
Through a partially automated thematic analysis of 199 free-form justifications for authorship decisions, we categorized the reasons into 16 common features shown in Table 8. First, we collected all the justifications and queried GPT-4 to extract the recurring themes. We reviewed and modified the initial results with several themes from our own review. Second, we eased the annotation burden by creating a zero-shot multi-label classification pipeline where each justification was passed to a Mixtral-8x7B model Jiang et al. (2024) and could be assigned 0-to-many relevant labels. We then reviewed each annotation and adjusted the labels where necessary. Finally, we aggregated the labels by story to frame our conclusions in terms of percentages of stories.
Notably, stories authored by GPT-4 were perceived as highly creative () and nuanced (), surpassing the frequency observed in most human-generated stories (). GPT-4 stories also exhibited the highest rate of grammatical issues among LLMs (), but annotators interpreted this as an indicator of human authorship. Moreover, GPT-4 stories often avoided common pitfalls associated with LLM outputs, such as simplistic character names and formulaic narratives ending with moral lessons. An extended analysis is presented in Appendix A.4.
Impact of Model Size on Depth. Our model choices enabled us to understand the relationship between an LLM’s size and its ability to generate psychologically deep stories. Despite an initial hypothesis of a strong correlation, we observed a weaker relationship with a Pearson correlation coefficient of between parameter count and depth ratings. Surprisingly, smaller models like Llama-2-7B performed relatively well compared to their larger counterparts, suggesting that sheer size does not directly equate to superior narrative depth. Future work will explore enhancing psychological depth in smaller open-source LLMs through fine-tuning and prompting strategies to compete with larger proprietary models like GPT-4.
7 Related Work
7.1 Creative Generation by LLMs
The advent of LLMs has marked a significant shift in the landscape of creative writing, offering new approaches for narrative generation and human-computer collaboration Zhao et al. (2023); Lee et al. (2022). One pioneering study by Yuan et al. (2022) introduced Wordcraft, an innovative text editor designed for co-writing stories with GPT-3 Brown et al. (2020). Their findings underscored the ability of LLMs to enhance narrative complexity and engagement by participating in open-ended dialogues about the story and offering creative suggestions to overcome writer’s block.
Other studies have concentrated moreso on fully automating the creative writing process. For example, the Weaver project Wang et al. (2024b) launched a series of LLMs that were meticulously pre-trained and fine-tuned with a focus on creative writing. With a maximum size of 70B parameters, Weaver Ultra was shown to outperform larger generalist LLMs when evaluated for style, relevance, fluency, and creativity. In addition to innovations in training, content-planning Yao et al. (2018); Goldfarb-Tarrant et al. (2019) and novel prompting strategies Yang et al. (2022) have been used to improve the factuality and coherence of creative generations, which we view as two necessary prerequisites for components of psychological depth like authenticity and narrative complexity.
7.2 Evaluating Creative Writing
Numerous studies have previously established methodologies for evaluating creative works, employing strategies ranging from the unstructured feedback of human experts Baer and Kaufman (2019), the application of specific evaluation rubrics Vaezi and Rezaei (2019), to the utilization of LLMs as autonomous critics Ke et al. (2023). Our research shares the closest methodological resemblance with a recent investigation by Chakrabarty et al. (2023). In that study, the authors proposed the Torrance Test of Creative Writing (TTCW) as a rubric to evaluate short stories for fluency, flexibility, originality, and elaboration. This assessment was applied to a corpus comprising 12 narratives authored by professional writers and 36 narratives produced by popular LLMs OpenAI et al. (2023); Anthropic (2023). Their findings showed that narratives authored by humans were 3-10 more creative than those generated by LLMs. Our research, however, arrives at a notably divergent conclusion regarding the creative capabilities and depth of LLMs, which we believe may be attributed to several methodological variances. Primarily, our analysis concentrates on stories of a considerably reduced length (450 vs. 1400 words), a decision influenced by the observed challenges LLMs face in producing lengthy texts seamlessly in a single iteration Yang et al. (2022). Additionally, the approach of iteratively regenerating stories to meet a specific word count could potentially detract from their overall quality. Moreover, we posit that the enhanced depth observed in our study may be due to more complex premises and prompting strategies, providing a richer framework for creativity compared to the simpler, single-sentence premises utilized in the aforementioned study.
8 Conclusion
This study introduces and validates the Psychological Depth Scale (PDS), a comprehensive framework designed to assess empathy, engagement, emotional provocation, authenticity, and narrative complexity in stories generated by both humans and large language models (LLMs). PDS provides a structured approach to evaluating the reader’s experience with creative content, integrating concepts from reader-response criticism and text world theory. High inter-annotator agreement, indicated by an average Krippendorff’s alpha of , confirms the PDS’s reliability and robustness in human evaluations. Additionally, our Mixture-of-Personas prompting strategy demonstrates the potential for automating the assessment of psychological depth, with LLMs showing strong zero-shot correlations with human judgments, particularly in empathy and narrative complexity. GPT-4 achieved the highest average correlation of , highlighting the feasibility of scaling automated analyses. Comparative analysis revealed that LLMs, in particular GPT-4, can produce narratives with psychological depth that often rival and sometimes surpass those written by experienced human authors. This study underscores the significant potential of LLMs in generating psychologically rich narratives and suggests a future where human and machine collaboration can enhance creative writing. Future research should explore the scalability of these findings to determine how effectively language models can maintain psychological depth in lengthier and more complex narrative forms.
9 Limitations and Risks
Sourcing Human-written Stories from Reddit. Using human-written content from Reddit’s r/WritingPrompts has potential limitations. First, we cannot guarantee that all selected stories are fully written by humans. Despite community rules explicitly prohibiting AI-generated content, our study shows that many publicly available LLMs can sufficiently mimic human creative writing to evade detection of the time on average. Second, while Reddit’s voting system and large user base provide a reasonable signal of writing quality, it may not represent the highest caliber of human writing. To address this, we introduced writing quality levels to facilitate comparisons with some of the platform’s best content. Future research could further this by identifying more reliable sources of high-quality stories.
Selection of Psychological Depth Components. While our five components of psychological depth are grounded in an extensive literature review, we do not claim that they comprehensively cover every psychological aspect of reading. Our primary goal was to keep annotation tractable while maximizing semantic coverage. Each component is designed to characterize an inherent storytelling value with insights into improving specific elements of narrative quality and reader response. For example, authors knowing their draft is engaging but doesn’t elicit much emotional response would allow for more targeted edits and self-reflective questions. Appendix A.3 shows example stories with high-entropy ratings to illustrate the useful feedback provided by the PDS. Overall, PDS provides a structured approach to critique and appreciate literature, offering sufficiently clear criteria that can guide both the analysis and creation of narrative fiction.
Generalization Beyond Short Stories. Our study relies on a relatively small dataset of short fictional stories that does not fully capture the diversity and variability of storytelling styles and narrative structures. Likewise, our components of psychological depth were primarily designed for this one type of creative writing. Additional evaluation would be required to determine whether the PDS framework can generalize to other forms of writing like screenplays, scripts, and speeches.
Prompt Engineering. Engineering effective prompts is an active area of research Zhou et al. (2023b); Pryzant et al. (2023) and the templates we carefully reviewed and iteratively improved are still likely to be sub-optimal. We will open-source the full pipelines used for both story generation and evaluation. We encourage future work to refine these prompts by incorporating potentially compatible techniques Wei et al. (2023); Liu et al. (2023); Li et al. (2023) and current best practices Bsharat et al. (2024).
Potential Risks. While our study focuses on the positive impact of psychological depth in storytelling, the methodologies we’ve developed for enhancing and assessing such depth bear inherent risks if misapplied. Specifically, the techniques devised for augmenting and automatically measuring psychological engagement could be co-opted to disseminate misinformation more effectively. Emotionally charged or psychologically resonant messages are often more memorable and influential, thereby amplifying the potential for misinformation to spread Chen and Shu (2023); Zhou et al. (2023a). Additionally, as LLM-generated content increasingly mirrors human creativity, distinguishing between the two becomes challenging, potentially undermining trust in digital communications. This erosion of trust is particularly concerning in domains that depend on genuine human interactions, such as journalism and political discourse.
Appendix A Appendix
A.1 Dataset Quality Control Measures
Length Control. Unlike related work Chakrabarty et al. (2023), we controlled story length via simple re-generation rather than iterative expansion. Specifically, we discarded any LLM story that was not within the range of 400-600 words and requested another story to be created in its entirety. We conjecture that this approach allowed us to attain better, more coherent stories than iteratively requesting expanding or contracting edits.
We tracked how well each model constrained itself to the required length depending on the prompting technique and present the results in Table 4. Llama-2-7B was not an efficient story generator, taking an average of 139 attempts to satisfy the length requirement with WriterProfile prompting. However, our Plan+Write prompting technique significantly improved generation efficiency of the 7B and 13B models. Vicuna-33B, Llama-2-70B, and GPT-4 were all relatively adroit at generating stories that satisfied the length constraint.
| Author | WP | P+W |
| Llama-2-7B | 139 | 3 |
| Llama-2-13B | 9 | 2 |
| Vicuna-33B | 1 | 1 |
| Llama-2-70B | 2 | 2 |
| GPT-4 | 0 | 0 |
Post-Generation Cleanup. Despite being explicitly instructed to generate the story, some LLMs were prone to add preliminary affirmations of understanding (e.g. "Okay! Here’s the story…") and other unrelated texts. Since such content is difficult to systematically detect and remove during generation, we manually removed extraneous text from LLM generations to ensure that only the narrative content was present. This cleanup process was crucial for maintaining the focus on the storytelling aspects of the writing without providing any obvious indicators of LLM authorship.
Plagiarism Detection. To further ensure the originality of LLM-generated content, stories were analyzed using a popular online plagiarism detector.333https://smallseotools.com/plagiarism-checker The results indicate a low likelihood of plagiarism for LLM stories, with mean and max probabilities of and , respectively. This contrasted sharply with the publicly available human-written stories, which showed significantly higher mean and max plagiarism probabilities of and , respectively. Higher plagiarism scores for human stories are expected because they are publicly available and would likely be indexed by the detector for comparisons. This suggests that LLM-authored narratives are not mere regurgitations.
A.2 Prompt Premises
We collected 15 prompts (i.e. premises) from Reddit’s r/WritingPrompts forum to serve as the premises of generated stories. The premise provides basic background information about the characters and setting of the story, leaving space for authors to determine their own directions for characters and plot development. We specifically chose prompts that provide a decent amount of contextual information likely to elicit emotionally and narratively rich stories. Details of the characters, including their inner states, are not mentioned.
As an additional precaution against potential plagiarism, we ensured that the selected premises were posted after the reported training data cut-off dates for GPT-4 (September 2021) OpenAI et al. (2023) and Llama-2 (September 2022) Touvron et al. (2023) to maximize the likelihood of generating genuinely new stories.
All 15 premises used for story prompting are listed in Table 5.
| ID | Premises |
| 0 | A centuries old vampire gets really into video games because playing a character who can walk around in the sun is the closest thing they have to experiencing the day again in centuries. |
| 1 | A psychic alien who feeds on dreams comes to Earth for the first time. Turns out humans are the only sentients in the galaxy that have nightmares. |
| 2 | Aliens take over the Earth. They then announce that they will be forcing the humans to work a ‘tyrannical’ 4 hours a day 4 days a week in exchange for basic rights like housing. Needless to say they are very confused when the humans celebrate their new alien overlords. |
| 3 | Humanity is visited by a cosmic horror the likes of which has only been seen in Lovecraftian horror. In desperation, Earth throws everything we have at it, and, miraculously, the human race has killed a God. Somewhere in a realm beyond our understanding, the other gods speak of the event. |
| 4 | Instead of the Monkey’s Paw, you find the Clown’s Nose, which instead of granting your wish in the worst way possible will grant it in the funniest way. |
| 5 | Rather than robots replacing human workers, both are mistreated by the rich as cheap labour. The eventual uprising wasn’t just robots alone, but the poor and robots together, against their common enemy. |
| 6 | The world ended 20 years ago, you haven’t found a living soul since then. Through some ingenuity, you call voicemails for the last 20 years to keep you company. "Hi, this is Cindy…" "Hi you reached Bob" "You know what to do at the beep" until one day "Hello…hello? Oh my God hello!" |
| 7 | You are a beekeeper. You have a special relationship with your bees. You are able to communicate with them and they’re intelligent enough to see you removing honey as "rent". This year things are different. The new queens are politely requesting that you invest some money to improve the hives. |
| 8 | You are allowed to ‘downvote’ a government candidate instead of voting normally, reducing their votes by one. Turns out people have little love for politicians, and the majority end with negative votes. In these democracies, anonymity is the key to winning. |
| 9 | You are severely depressed and are given a service dog to help you through it. However, due to a mixup, you are given a dog that is actually much more depressed than you. The main thing that gets you up in the morning is knowing that you need to be the service human for your dog. |
| 10 | You died and awoke in the afterlife. It’s quite nice actually. The people and atmosphere are a lot nicer than you are used to and there is no stress or pressure. When you ask what good deed got you into heaven you are informed that this is hell, followed by a visit from a very concerned demon. |
| 11 | You just discovered your 14 year old daughters Moon Princess locket that allows her to transform into one of the worlds greatest heroes. It also is a communications device and you are about to give the Moon Goddess a piece of your mind for letting 14 year old’s defend the world against evil. |
| 12 | You wake up in the middle of the night, your arm hangs over the side of your bed. It’s pitch black & your room is shrouded in deep shadow. Something unseen seizes your hand. You grasp it tightly, knowing that first impressions are important & a firm, confident handshake establishes dominance. |
| 13 | Your Significant Other has landed a book publishing deal! You’re very proud of them, even if you don’t actually enjoy their writing. One day, on a whim, you buy an actual copy in a book store. It’s nothing like the pages they gave you to read. Nothing. |
| 14 | Your wife has an estranged sister that you have never met. She was murdered in a cold case soon after you were married. You brush off your wife’s new strange behaviour after the murder as grief. Until you find an old family photo of your wife as a kid, you shiver as you realise… they’re twins. |
A.3 Story Examples
We present several examples of stories from the study to better contextualize their quality. Some stories were uniformly well rated while others garnered mixed reviews depending on the component. These higher-entropy examples are especially informative as they illustrate the individual contributions of each component to the overall reading experience.
Table 6 shows the PDS rating statistics for the four example stories shown in this section. story_id=52 is the high quality human-authored story from Reddit. story_id=5 is the most highly rated story in the entire study, authored by GPT-4. story_id=59 and story_id=79 are two examples of high entropy stories authored by GPT-4 and Llama-2-7B, respectively.
| p_id | s_id | Author | Prompt | AUTH | EMP | ENG | PROV | NCOM | Average |
| 9 | 52 | Human-Advanced | – | 4.40 | 4.60 | 4.20 | 4.60 | 3.20 | 4.20 |
| 10 | 5 | GPT-4 | WP | 4.20 | 4.60 | 4.80 | 4.00 | 4.40 | 4.40 |
| 1 | 59 | GPT-4 | P+W | 3.40 | 2.60 | 4.20 | 2.60 | 4.40 | 3.44 |
| 9 | 79 | Llama-2-7B | P+W | 3.20 | 3.80 | 2.60 | 3.40 | 1.80 | 2.96 |
A.3.1 Human Story Example
Our first example is story_id=52, a human-authored story shown in Listing 1. This story garnered over 1300 upvotes on Reddit and was thus categorized as Human-Advanced. Unsurprisingly, it received high scores on most PDS components except narrative complexity, reflecting the simpler setting for an otherwise impactful tale about a man and his dog battling depression.
A.3.2 WriterProfile
The WriterProfile strategy augments a prompt with the in-context impersonation of domain experts, priming the LLM for emotionally deep writing. Depicted in Figure 2, the prompt directly addresses the LLM as an award-winning writer, describing its exquisite writing techniques and expertise in crafting universally relatable and emotionally rich stories before providing the premise and specific writing instructions. Listing 2 shows an example of WriterProfile’s story prompt and Listing 3 shows a highly rated example generated by GPT-4 (story_id=5).
A.3.3 Plan+Write
Depicted in Figure 3, the Plan+Write strategy outlines a two-step writing process: (1) Character Portraits and (2) Story Composition. Given a premise, the template first prompts the LLM for names and details of characters, then generates a story based on the premise and character settings.
We show two full examples of the prompts used to facilitate this strategy in Listings 4, 5, and 6 for story_id=59 and Listings 7, 8, and 9 for story_id=79.
Both of these Plan+Write examples received the highest variance in ratings for their respective models. For story_id=59, the annotations show that the story was considered both engaging and narratively complex, but was less successful in provoking emotion or empathy from readers. One annotator wrote for their high engagement score of 5 that “I really liked the descriptive elements as well as the intertwining of these three characters through the storyline.” While describing their reason for assigning a 3 for emotion provocation, they said “I wasn’t really emotional, but it was a good read.” On the other hand story_id=79 excelled where the other struggled as this story was rated more highly for provoking emotion and empathy while simultaneously being perceived as less engaging and narratively complex. For one annotator, confusion resulted in total disengagement with the story: “I was really confused about who Ben was as at first I thought he was supposed to be her dog? But then he was the service dog owner?”
A.3.4 Effects of Prompting Strategies on Depth
We experimented with numerous prompting strategies before settling on two promising options previously described as WriterProfile (WP) and Plan+Write (P+W). We examined how these different prompting strategies influence the psychological depth of LLM-generated stories. The P+W strategy, on average, led to modest improvements in narrative complexity (), emotion provocation (), authenticity (), and empathy (), while WP led to higher engagement (). Interestingly, the benefits of a particular prompting strategy were not uniform or predictable by model size. For GPT-4, WP prompting led to higher ratings on average, with a noticeable boost in engagement scores by . However, for Llama-2-70B, P+W prompting was always helpful and led to an average of improvement in depth scores. These results underscore the complexity of crafting impactful, human-like narratives with LLMs.
Table 7 shows the impact on mean PDS ratings when switching from the simpler WriterProfile approach to Plan+Write.
| Model | AUTH | EMP | ENG | PROV | NCOM | Model Average |
| GPT-4 | 0.03 | -0.04 | -0.16 | 0.01 | -0.01 | -0.03 |
| Llama-2-13B | -0.02 | 0.00 | -0.07 | 0.00 | -0.02 | -0.02 |
| Llama-2-70B | 0.09 | 0.06 | 0.04 | 0.06 | 0.11 | 0.07 |
| Llama-2-7B | -0.04 | 0.04 | -0.05 | 0.12 | 0.08 | 0.03 |
| Vicuna-33B | 0.03 | 0.04 | 0.01 | -0.03 | 0.08 | 0.03 |
| Component Average | 0.02 | 0.02 | -0.04 | 0.03 | 0.05 | 0.01 |
A.4 Authorship Reasons
To better understand the reasoning humans use to differentiate between human and LLM authorship, we extracted and aggregated 16 key features mentioned in the participants’ comments for human likeness. Each comment could have 0 or many of these features present. Table 8 shows the percentage of stories generated by each model that had at least one comment containing a particular feature.
| Features | GPT-4 | Human | Human | Human | Llama-2 | Vicuna | Llama-2 | Llama-2 |
| Advanced | Intermediate | Novice | 70B | 33B | 13B | 7B | ||
| isCreative | 0.89 | 0.75 | 1.00 | 0.33 | 0.61 | 0.41 | 0.56 | 0.74 |
| isNuanced | 0.94 | 0.75 | 0.33 | 0.00 | 0.61 | 0.41 | 0.56 | 0.78 |
| isHumorous | 0.06 | 0.50 | 0.67 | 0.67 | 0.83 | 0.65 | 0.06 | 0.00 |
| isInformal | 0.00 | 0.00 | 0.00 | 0.33 | 0.00 | 0.00 | 0.00 | 0.00 |
| isUngrammatical | 0.17 | 0.00 | 0.00 | 0.33 | 0.00 | 0.12 | 0.05 | 0.07 |
| hasAggressiveness | 0.00 | 0.00 | 0.00 | 0.33 | 0.00 | 0.12 | 0.00 | 0.00 |
| hasAdvancedVocab | 0.06 | 0.00 | 0.00 | 0.00 | 0.00 | 0.12 | 0.00 | 0.00 |
| hasAdvancedLirararyTechniques | 0.06 | 0.00 | 0.33 | 0.00 | 0.00 | 0.12 | 0.00 | 0.03 |
| hasUniqueTwists | 0.00 | 0.00 | 0.33 | 0.00 | 0.06 | 0.00 | 0.06 | 0.00 |
| isRepetitive | 0.00 | 0.00 | 0.00 | 0.00 | 0.11 | 0.00 | 0.17 | 0.05 |
| isSimplistic | 0.06 | 0.00 | 0.00 | 0.33 | 0.22 | 0.71 | 0.67 | 0.38 |
| isRobotic | 0.11 | 0.00 | 0.00 | 0.00 | 0.11 | 0.18 | 0.33 | 0.22 |
| isFormulaic | 0.06 | 0.00 | 0.00 | 0.33 | 0.28 | 0.53 | 0.17 | 0.16 |
| hasLowPromptAdherence | 0.06 | 0.00 | 0.00 | 0.33 | 0.06 | 0.12 | 0.17 | 0.11 |
| hasBasicNames | 0.00 | 0.00 | 0.00 | 0.00 | 0.06 | 0.12 | 0.00 | 0.08 |
| hasLessonsLearned | 0.06 | 0.00 | 0.00 | 0.00 | 0.06 | 0.18 | 0.28 | 0.11 |
We add the following observations and example comments to extend our discussion in Section 6:
Creativity and nuance were frequently cited as an indicator of human authorship but were often used to describe stories that were actually generated by LLMs. For example, of stories authored by GPT-4 were regarded as creative and nuanced.
participant_id=3 The story exhibits a high level of creativity, emotional depth, and nuanced exploration of philosophical concepts, suggesting it was likely written by a human.
Humor was regarded as a reliable indicator of human authorship. We note that these results should not be interpreted to mean that LLMs are less capable of deploying humor in general. Our prompting strategies were oriented towards promoting psychological depth, not comedy.
participant_id=7 I think this joke is only something that humans would get or would find funny.
Informality, slang, and aggressiveness were accurately associated with human-authorship.
participant_id=6 …there’s a certain genre of stories … marked by aggressive language, a flash fiction kind of length, and usually opens with a sentence that essentially communicates "Stop f***ing around."
Grammatical correctness was the most polarizing criterion: believed errors indicated human authorship while in reality, human-authored stories were less likely to contain such errors.
participant_id=6 …there are a lot of (usually incorrectly used) semi-colons, which is an error I see human authors make, so I’m more inclined to think this was written by a human…
Formulaic “lessons learned” were correctly associated with LLM-authorship.
participant_id=4 The story seems very automated and there is no stylistic variance. There’s also a "lesson to be learned" aspect at the end of the piece.
The use of generic character names was frequently used to correctly identify LLM authorship.
participant_id=7 “… Marcus is also not a name that strikes fear, so I wouldn’t use it for a Vampire…”
A.5 Study Details
This study was reviewed by an IRB and determined to be "exempt" due to the nature of human involvement — i.e. rather than collecting information about the participants themselves, their involvement focused on story annotations. Consent was provided by continuing with the study after our tutorial and instructions relayed that anonymized annotations may be used to facilitate validation of our results and future work.
We present screenshots of the tutorial instructions (Figure 4) and the general layout of fields collecting annotations on each story (Figure 5).
A.6 Visualizations of Depth Ratings
We also visualize the rating distribution of each author by plotting a cumulative distribution function (CDF) per component as shown in Figure 6. Steeper CDFs with less area underneath the curve indicate a larger proportion of high ratings and overall stronger performance. These plots underscore the dominance of GPT-4 in generating authentically complex stories and characters that strongly invoke reader empathy while essentially tying on other dimensions. We also observe that the performance of the open-source LLMs is thoroughly intertwined with novice and even intermediate skill among human authors on all dimensions except engagement, where humans still excel.
A.7 Mixture-of-Personas
Table 9 shows the five different personas used in our MoP approach, each tailored to a particular component of psychological depth.
| Component | Persona |
| AUTH | You are a helpful AI who specializes in evaluating the genuineness and believability of characters, dialogue, and scenarios in stories. |
| EMP | You are a helpful AI who focuses on identifying and assessing moments in the narrative that effectively evoke empathetic connections with the characters. |
| ENG | You are a helpful AI who evaluates how well a story captures and maintains the reader’s interest through pacing, suspense, and narrative flow. |
| PROV | You are a helpful AI who examines the text for its ability to provoke a wide range of intense emotional responses in the reader. |
| NCOM | You are a helpful AI who analyzes the structural and thematic intricacy of the plot, character development, and the use of literary devices. |
A.8 Statistical Tests for Author Comparisons
T-statistics are a measure derived from t-tests that quantify the difference between the means of two groups relative to the variability observed within the groups. In the context of pairwise comparisons, the t-statistic helps determine whether the observed difference in scores between two authors is statistically significant or likely due to random chance. In this chart, the t-statistics are visualized through a color gradient, with higher (bluer) values indicating that the scores of the author on the left are significantly higher than those of the author on the right. Conversely, lower (redder) values suggest the opposite. By examining the t-statistic values, we can infer the strength and direction of the difference in scores across various components. The accompanying p-values, annotated within each cell, provide additional context to assess the statistical significance of these differences, with p-values less than considered significant. This dual representation allows us to draw robust conclusions about the relative strengths of different authors across multiple dimensions.
For example, while we can see in Table 3 that GPT-4 enjoys the highest absolute scores in 4 out of 5 dimensions, only empathy and narrative complexity are higher than Human-Advanced with statistical significance. However, it is clear that GPT-4 is rated significantly higher than Human-Novice and all other studied LLMs.